# Diagonalization $of$ Polynomial-Time Deterministic Turing Machines via Nondeterministic Turing Machines

### v31

@@
Abstract
The diagonalization technique was invented by Georg Cantor to show that there are more real numbers than algebraic numbers and is very important in theoretical computer science. In this work, we enumerate all of the polynomial-time deterministic Turing machines and diagonalize against all of them by a universal nondeterministic Turing machine. As a result, we obtain that there is a language $L_{d}$ not accepted by any polynomial-time deterministic Turing machines but accepted by a nondeterministic Turing machine running within time $O(n^{k})$ for any $k\in\mathbb{N}_{1}$. Based on these, we further show that $L_{d}\in\mathcal{NP}$ That is, we present a proof that $\mathcal{P}$ and $\mathcal{NP}$ differ. Meanwhile, we show that there exists a language $L_{s}$ in $\mathcal{P}$ but the machine accepting it also runs within time $O(n^{k})$ for all $k\in\mathbb{N}_{1}$. Lastly, we show that if $\mathcal{P}^{O}=\mathcal{NP}^{O}$ and on some rational base assumptions then the set $P^{O}$ of all polynomial-time deterministic oracle Turing machines with oracle $O$ is not enumerable, thus demonstrating that diagonalization technique (via a universal nondeterministic oracle Turing machine) will generally not apply to the relativized versions of the $\mathcal{P}$ versus $\mathcal{NP}$ problem.
@@ 1. Introduction
2. Preliminaries
5. Proof of $L_{d}\in\mathcal{NP}$
7. Is “Relativization Barrier” Really A Barrier
8. Conclusions
Appendix A. The Two Definitions of $\mathcal{P}$ (Respectively, $\mathcal{NP}$) Are Equivalent
1 Introduction
@@
In $1936$, Turing’s landmark $36$-page paper [Tur37] opened the door to the computer science, which has since evolved into numerous subfields such as computability theory, formal language and automata theory, computational complexity theory, algorithm theory, and so on. Turing’s contributions, on the one hand, were highly influential in the development of theoretical computer science that he is widely regarded as the father of theoretical computer science [A4]. But on the other hand, although Turing’s work initiated the study of theoretical computer science, he was not concerned with the efficiency of his machines, which is the main topic in computational complexity theory. In fact, Turing’s concern [Tur37] was whether they can simulate arbitrary algorithms given sufficient time (see e.g. [Coo00]).
@@
The computational complexity theory is a central subfield of the theoretical foundations of computer science, which mainly concerns the efficiency of Turing machines (algorithms), or the intrinsic complexity of computational tasks, i.e., focuses on classifying computational problems according to their resource usage, and relating these classes to each other (see e.g. [A5]). In other words, it specifically deals with fundamental questions such as what is feasible computation and what can and cannot be computed with a reasonable amount of computational resources in terms of time or space. In short, the theory formalizes this intuition to study these problems and quantifying their computational complexity.
@@
There are different measures of difficulty to the computing such as the amount of communication, the most basic and fundamental of which that appear particularly important are perhaps time and space. The fundamental measure of time opened the door to the study of the extremely expressive time complexity class $\mathcal{NP}$, one of the most important classical complexity classes, i.e., nondeterministic polynomial-time. This class comprises languages that can be computed in polynomial-time by nondeterministic Turing machines. The famous Cook-Levin theorem [Coo71, Lev73] shows that this class has complete problems, which states that the Satisfiability is $\mathcal{NP}$-complete, i.e., Satisfiability is in $\mathcal{NP}$ and any other language in $\mathcal{NP}$ can be reduced to it in polynomial-time. This famous result also opened the door to research into the rich theory of $\mathcal{NP}$-completeness [Kar72].
@@
The famous $\mathcal{P}$ versus $\mathcal{NP}$ problem, which unquestionably has caught the interest of the mathematical community (see e.g. [Coo03, BCSS98, Sma00]), is a major open question in theoretical computer science community, specifically, in computational complexity theory community. The question asks whether every problem in $\mathcal{NP}$ can also be solved in polynomial-time by a deterministic Turing machine. It appeared explicitly for the first time in the papers of Cook [Coo71], Karp [Kar72] and Levin [Lev73]. From the point of view in [Bus12], one reason for why this problem catching the interest of the mathematical community [BCSS98, Sma00, Coo03] is that $\mathcal{P}=\mathcal{NP}$ could make the practice of mathematics too easy. Mathematical research could be automated by formalizing mathematical questions completely, and then blindly searching for proofs of conjectured mathematical statements. If $\mathcal{P}=\mathcal{NP}$, this process could succeed whenever proofs are not to large. This would obviously be a major change in the practice of mathematics! It is also worth focusing on that, the relationships between the notion of one-way functions (OWFs) [DH76] in modern cryptography (i.e., one-way functions in the average-case model) and the $\mathcal{P}$ versus $\mathcal{NP}$ problem are interesting topics and are discussed in a standard cryptographic textbook [Gol01] by Goldreich. In addition, the relationships between one-way functions in the worst-case model and the $\mathcal{P}$ versus $\mathcal{NP}$ problem are also discussed in [Pap94] (see p. 281–284 in [Pap94], or [Ko85, GS88]). For more other details, the reader can find that the reference [Wig07] also contains a detailed background about it.
@@
There are serval different and equivalent formulations of the $\mathcal{P}$ versus $\mathcal{NP}$ problem, the interested reader is referred to [May04]. In particular, the logical characterizations of the question “is $\mathcal{P}$ a proper subset of $\mathcal{NP}$” can be reformulated as “is existential second-order logic able to describe languages (of finite linearly ordered structures with nontrivial signature) that first-order logic with least fixed point cannot?”; see e.g. [A1, May04].
@@
Figure 1: Euler diagram for $\mathcal{P}$, $\mathcal{NP}$, $\mathcal{NP}$-complete and $\mathcal{NP}$-hard set of problems
@@
To the best of our knowledge, many of the methods used to attack the $\mathcal{P}$ versus $\mathcal{NP}$ problem (in the direction of $\mathcal{P}\neq\mathcal{NP}$) have been combinatorial or algebraic; for example, circuit lower bounds, diagonalization, and relativization [BGS75], but previously all of these attempts fail (see e.g. [Coo00, Wig07, For09]). In 1997, Razborov and Rudich defined a general class of proof techniques for circuit complexity lower bounds (see [Raz85]), called natural proofs [RR97]. At the time, all previously known circuit lower bounds were natural, and circuit complexity was considered a very promising approach for resolving $\mathcal{P}$ versus $\mathcal{NP}$ problem. However, in [RR97], Razborov and Rudich showed that if one-way functions exist, then no natural proof method can distinguish between $\mathcal{P}$ and $\mathcal{NP}$ (see e.g. [A1]). Although one-way functions have never been formally proven to exist, most mathematicians believe that they do, and a proof of their existence would be a much stronger statement than $\mathcal{P}\neq\mathcal{NP}$. Thus it is unlikely that natural proofs alone can resolve the $\mathcal{P}$ versus $\mathcal{NP}$ problem (see e.g. [A1]). But, maybe some day in the future, someone is going to present a proof of that $\mathcal{NP}\not\subseteq\mathcal{P}/poly$ by circuit lower bounds (for example, to prove $3$-SAT or other $\mathcal{NP}$-complete problems not in $\mathcal{P}/ploy$ by circuit lower bounds). In view of this, we believe that only the God knows whether the circuit lower bounds has really a natural proofs barrier to the $\mathcal{P}$ versus $\mathcal{NP}$ problem.
@@
More interestingly, the technique of diagonalization led to some early successes in computational complexity theory. In history, it was invented by Georg Cantor [Can91] to show that there are more real numbers than algebraic numbers, then it was used by Alan Turing to show that there exists sequence which is not computable (see [Tur37], p. 246), and then refined to prove computational complexity lower bounds by Hartmanis and Stearns [HS65]. A typical theorem in this area is that more time/space buys more computational power [HS65, Coo73, FS07, Pap94, Sip13, AB09]. For instance, there are functions computable in time $n^{2}$, say, which are not computable in time $n$. At present, we assume that the heart of such arguments is the existence of a universal Turing machine, which can simulate every other Turing machine with only a small loss in efficiency.111 At the end of the work, it will become clear that the notion of Enumerable of Turing machines is an important prerequisite for application of diagonalization.
@@
As we mentioned earlier, the diagonalization technique was invented by Cantor [Can91] to show that there are more real numbers than algebraic numbers and early it demonstrated powers in separating two complexity classes. There is an important doubt: can diagonalization approach resolve the $\mathcal{P}$ versus $\mathcal{NP}$ problem? Researchers realized in the $1970$s that diagonalization alone may not be able to resolve the $\mathcal{P}$ versus $\mathcal{NP}$ problem; it needs other new techniques besides diagonalization (see e.g. [Sip13, AB09]). But, in the end of the work, we will say that the above point of view is partially correct and leave aside why we think so.
@@
It is interesting that the complexity class $\mathcal{NP}$ has a rich structure under the assumption that $\mathcal{P}$ and $\mathcal{NP}$ differ. In [Lad75], Lander constructed a language which is $\mathcal{NP}$-intermediate by the method of lazy diagonalization under the assumption that $\mathcal{P}\neq\mathcal{NP}$ (there are other problems which are suspected of being $\mathcal{NP}$-intermediate, for example, the graph isomorphism problem believed to be at least not $\mathcal{NP}$-complete [AK06]). We noted that the lazy diagonalization put the $\mathcal{NP}$-intermediate language accepted by no polynomial-time deterministic Turing machine, and our curious here, after reading [Lad75] for many times, is that why Lander did not diagonalize against all polynomial-time deterministic Turing machines directly to see whether it is possible to give a solution of the $\mathcal{P}$ versus $\mathcal{NP}$ problem? Because it is a widespread belief that $\mathcal{P}$ and $\mathcal{NP}$ are different, we naturally would consider whether we can diagonalize against all of the polynomial-time deterministic Turing machines by a universal nondeterministic Turing machine to produce a language accepted by no polynomial-time deterministic Turing machine but accepted by some nondeterministic Turing machine. Once we have constructed a language that is not in $\mathcal{P}$ by diagonalization, we can do the second half of the work to see whether this language is in $\mathcal{NP}$ or not.
@@
Before introducing the main results, let us stress that the definitions about almost all of the formal concepts involved in the subsection below will be postponed and, due to our style of writing, we just suppose that the reader is familiar with these notions. To those who are not familiar with these concepts, we suggest to proceed to read the relevant definitions in Section 2 and in Section 7 first.
1.1 Main Results
@@
We put our aforementioned curious, or ideas into practice in this work. We first enumerate all polynomial-time deterministic Turing machines222 It is much suitable to say that we first present a way to enumerate all of the polynomial-time deterministic Turing machines. and then diagonalize against all of them by a universal nondeterministic Turing machine. Generally, complexity theorists used universal deterministic Turing machines to diagonalize against a list of deterministic Turing machines (see e.g. [HS65, Pap94, AB09]), or applied universal nondeterministic Turing machines to diagonalize against a list of nondeterministic Turing machines; see e.g. [AB09, For00, FS07, Zak83]. Using a universal nondeterministic Turing machine to diagonalize against a collection of deterministic Turing machines appears less often, except in the author’s recent work [Lin21a], and it is maybe a new attempt that may lead to success. As an amazing result, we obtain the following important theorem:
Theorem 1
@@
There exists a language $L_{d}$ which is not accepted by any polynomial-time deterministic Turing machines, but accepted by a nondeterministic Turing machine. Furthermore, this nondeterministic Turing machine runs within time $O(n^{k})$ for any $k\in\mathbb{N}_{1}$. By this, it can be shown that $L_{d}\in\mathcal{NP}$.
from which it immediately follows that
Corollary 2
@@
In order to contrast with the construction in the proof of Theorem 1, we also design a universal deterministic Turing machine $M^{\prime}_{0}$ that accepts all languages in $\mathcal{P}$, and this machine also runs within time $O(n^{k})$ for all $k\in\mathbb{N}_{1}$. In contrast, the deterministic universal Turing machine $M^{\prime}_{0}$ does not perform any diagonalization operation. More interestingly, we conjecture that the language accepted by this deterministic universal Turing machine $M^{\prime}_{0}$ cannot be accepted by any deterministic $O(n^{i})$ time-bounded Turing machine for fixed $i\in\mathbb{N}_{1}$. Thus, the following theorem is an interesting by-product of our similar but different construction in proof of the Theorem 1:
Theorem 4
@@
Now, let us turn to the so-titled “Relativization Barrier.”
@@
In $1975$, Baker, Gill, and Solovay [BGS75] presented a proof of that:333 The proof is via $\mathcal{PSPACE}=\mathcal{NPSPACE}$, whose proof is by the fact that the space is re-useable but not valid for time, obviously.
@@
What are the implications of the above relativized result for the corresponding unrelativized question? A number of interesting ideas have been proposed, as observed by Ko (see [Ko85]): for example, Baker, Gill and Solovay [BGS75] suggested that their result imply that ordinary diagonalization techniques are not capable of proving $\mathcal{P}\neq\mathcal{NP}$ (similar perspectives also followed in [Sip13, BC94]). However, Kozen [Koz78] disagreed with this point of view. Hartmanis and Hopcroft [HH76] pointed out the possibility of the axiomatic independence of the question $\mathcal{P}\overset{?}{=}\mathcal{NP}$. Bennett and Gill [BG81] showed that $\mathcal{P}^{B}=\mathcal{R}^{B}$ but $\mathcal{P}^{B}\neq\mathcal{NP}^{B}$ relative to a random oracle $B$. Based on their results, they proposed the random oracle hypothesis: for an “acceptable” relativized statement $S^{A}$, $S^{\emptyset}$ is true if and only if $S^{A}$ is true with probability $1$ when $A$ is random.
@@
To this day, almost all complexity theory experts notice that the proof technique used to prove $\mathcal{P}\neq EXP$, i.e., diagonalization, would also ‘apply verbatim’ if we added an arbitrary oracle $O$. Thus, for any oracle $O$, we have $\mathcal{P}^{O}\neq EXP^{O}$. However, if we used similar techniques to show that $\mathcal{P}\neq\mathcal{NP}$, then it would also follow that $\mathcal{P}^{O}\neq\mathcal{NP}^{O}$ for all oracle $O$, which contradicts the result of [BGS75] above. This is the so-called “Relativization Barrier,” which almost all complexity theory experts think that any proof technique leading to $\mathcal{P}\neq\mathcal{NP}$ should overcome. Simultaneously, this is also the most notable difficulty on the road to attack this problem in the direction of $\mathcal{P}\neq\mathcal{NP}$ prior to our work. Based on the original perspectives given in [BGS75], it seems that the motivation of [BGS75] is to prove that $\mathcal{P}\neq\mathcal{NP}$ implies that $\mathcal{P}^{O}\neq\mathcal{NP}^{O}$ for all oracle $O$, via relativize, i.e., a proof technique invariant to adding oracles, but nevertheless we prefer to regard that $\mathcal{P}\neq\mathcal{NP}$ is not necessarily a necessary and sufficient condition for $\mathcal{P}^{O}\neq\mathcal{NP}^{O}$ for all oracle $O$444 For example, in the case where the set of $P^{O}$ is not enumerable, as will be explained below in detail.
@@
On the other hand, the role of relativization in complexity theory is interesting and important, and has been a central theme in complexity theory for almost two decades. Following the work [BGS75], much effort was expended to find contradictory relativizations for other open problems in complexity theory. A major part of this effort revolved around obtaining relativized results about the polynomial time hierarchy and its relationship to other classes; see e.g. [HCCRR93]. For example, contradictory relativizations for various unsolved problems regarding the polynomial hierarchy were found (see e.g. [HCCRR93]). These and other results led to a strong belief that problems with contradictory relativization are very hard to solve and are not amenable to current proof techniques, i.e., the solutions of such problems are beyond the current state of mathematics (see e.g. [Hop84, HCCRR93]).
@@
Now let us return to the proof technique of diagonalization once again. Cantor’s diagonal process, also called the diagonalization argument, was published in $1891$ by Georg Cantor [Can91] as a mathematical proof that there are infinite sets which cannot be put into one-to-one correspondence with the infinite set of positive numbers, i.e., $\mathbb{N}_{1}$ defined in the following Section 2. The technique of diagonalization was first used in computability theory in the $1930$’s by Turing [Tur37] to show that there exists undecidable language. In computational complexity theory, in their seminal paper [HS65], Hartmanis and Stearns employed the diagonalization proof to give time hierarchy. For more other summarization about it, please consult [For00] for a survey. On the other hand, some complexity theorists or textbooks (see e.g. page $73$ in [AB09]) in computational complexity theory regard that “diagonalization” is any technique that relies solely upon the following properties of Turing machines:
I

### v33

@@
Abstract.
The diagonalization technique was invented by Georg Cantor to show that there are more real numbers than algebraic numbers and is very important in theoretical computer science. In this work, we enumerate all of the polynomial-time deterministic Turing machines and diagonalize against all of them by a universal nondeterministic Turing machine. As a result, we obtain that there is a language $L_{d}$ not accepted by any polynomial-time deterministic Turing machines but accepted by a nondeterministic Turing machine running within time $O(n^{k})$ for any $k\in\mathbb{N}_{1}$. Based on these, we further show that $L_{d}\in\mathcal{NP}$. That is, we present a proof that $\mathcal{P}$ and $\mathcal{NP}$ differ. Meanwhile, we show that there exists a language $L_{s}$ in $\mathcal{P}$, but the machine accepting it also runs within time $O(n^{k})$ for all $k\in\mathbb{N}_{1}$. Lastly, we show that if $\mathcal{P}^{O}=\mathcal{NP}^{O}$ and on some rational base assumptions, then the set $P^{O}$ of all polynomial-time deterministic oracle Turing machines with oracle $O$ is not enumerable, thus demonstrating that the diagonalization technique (via a universal nondeterministic oracle Turing machine) will generally not apply to the relativized versions of the $\mathcal{P}$ versus $\mathcal{NP}$ problem.
keywords:
Diagonalization
keywords:
Polynomial-time deterministic Turing machine
keywords:
Universal nondeterministic Turing machines
@@ 1. Introduction
In $1936$, Turing’s landmark $36$-page paper [Tur37] opened the door to computer science, which has since evolved into numerous subfields such as computability theory, formal language and automata theory, computational complexity theory, algorithm theory, and so on. Turing’s contributions, on the one hand, were so highly influential in the development of theoretical computer science that he is widely regarded as the father of theoretical computer science [A4]. But on the other hand, although Turing’s work initiated the study of theoretical computer science, he was not concerned with the efficiency of his machines, which is the main topic in computational complexity theory. In fact, Turing’s concern [Tur37] was whether they can simulate arbitrary algorithms given sufficient time (see e.g., [Coo00]).
@@
The computational complexity theory is a central subfield of the theoretical foundations of computer science, which mainly concerns the efficiency of Turing machines (algorithms) or the intrinsic complexity of computational tasks, i.e., focuses on classifying computational problems according to their resource usage and relating these classes to each other (see e.g., [A5]). In other words, it specifically deals with fundamental questions such as what is feasible computation and what can and cannot be computed with a reasonable amount of computational resources in terms of time or space. In short, the theory formalizes this intuition to study these problems and quantify their computational complexity.
@@
There are different measures of difficulty to the computing, such as the amount of communication, the most basic and fundamental of which that appear particularly important are perhaps time and space. The fundamental measure of time opened the door to the study of the extremely expressive time complexity class $\mathcal{NP}$, one of the most important classical complexity classes, i.e., nondeterministic polynomial-time. This class comprises languages that can be computed in polynomial time by nondeterministic Turing machines. The famous Cook-Levin theorem [Coo71, Lev73] shows that this class has complete problems, which states that the Satisfiability is $\mathcal{NP}$-complete, i.e., Satisfiability is in $\mathcal{NP}$ and any other language in $\mathcal{NP}$ can be reduced to it in polynomial time. This famous result also opened the door to research into the rich theory of $\mathcal{NP}$-completeness [Kar72].
@@
The famous $\mathcal{P}$ versus $\mathcal{NP}$ problem, which unquestionably has caught the interest of the mathematical community (see e.g., [Coo03, BCSS98, Sma00]), is a major open question in the theoretical computer science community, specifically, in the computational complexity theory community. The question asks whether every problem in $\mathcal{NP}$ can also be solved in polynomial time by a deterministic Turing machine. It appeared explicitly for the first time in the papers of Cook [Coo71], Karp [Kar72], and Levin [Lev73]. From the point of view in [Bus12], one reason why this problem is catching the interest of the mathematical community [BCSS98, Sma00, Coo03] is that $\mathcal{P}=\mathcal{NP}$ could make the practice of mathematics too easy. Since mathematical research could be automated by formalizing mathematical questions completely and then blindly searching for proofs of conjectured mathematical statements. If $\mathcal{P}=\mathcal{NP}$, this process could succeed whenever proofs are not too large. This would obviously be a major change in the practice of mathematics! It is also worth focusing on that the relationships between the notion of one-way functions (OWFs) [DH76] in modern cryptography (i.e., one-way functions in the average-case model) and the $\mathcal{P}$ versus $\mathcal{NP}$ problem are interesting topics and are discussed in a standard cryptographic textbook [Gol01] by Goldreich. In addition, the relationships between one-way functions in the worst-case model and the $\mathcal{P}$ versus $\mathcal{NP}$ problem are also discussed in [Pap94] (see p. 281–284 in [Pap94], or [Ko85, GS88]). For more other details, the reader can find that the reference [Wig07] also contains a detailed background about it.
@@
There are several different and equivalent formulations of the $\mathcal{P}$ versus $\mathcal{NP}$ problem; the interested reader is referred to [May04]. In particular, the logical characterizations of the question “is $\mathcal{P}$ a proper subset of $\mathcal{NP}$” can be reformulated as “is existential second-order logic able to describe languages (of finite linearly ordered structures with nontrivial signature) that first-order logic with least fixed point cannot?”; see e.g., [A1, May04].
@@
Figure 1. Euler diagram for $\mathcal{P}$, $\mathcal{NP}$, $\mathcal{NP}$-complete, and $\mathcal{NP}$-hard sets of problems
@@
To the best of our knowledge, many of the methods used to attack the $\mathcal{P}$ versus $\mathcal{NP}$ problem (in the direction of $\mathcal{P}\neq\mathcal{NP}$) have been combinatorial or algebraic; for example, circuit lower bounds, diagonalization, and relativization [BGS75], but previously all of these attempts failed (see e.g., [Coo00, Wig07, For09]). In 1997, Razborov and Rudich defined a general class of proof techniques for circuit complexity lower bounds (see [Raz85]), called natural proofs [RR97]. At the time, all previously known circuit lower bounds were natural, and circuit complexity was considered a very promising approach for resolving the $\mathcal{P}$ versus $\mathcal{NP}$ problem. However, in [RR97], Razborov and Rudich showed that if one-way functions exist, then no natural proof method can distinguish between $\mathcal{P}$ and $\mathcal{NP}$ (see e.g., [A1]). Although one-way functions have never been formally proven to exist, most mathematicians believe that they do, and a proof of their existence would be a much stronger statement than $\mathcal{P}\neq\mathcal{NP}$. Thus it is unlikely that natural proofs alone can resolve the $\mathcal{P}$ versus $\mathcal{NP}$ problem (see e.g., [A1]). But, maybe some day in the future, someone is going to present a proof of that $\mathcal{NP}\not\subseteq\mathcal{P}/poly$ by circuit lower bounds (for example, to prove $3$-SAT or other $\mathcal{NP}$-complete problems not in $\mathcal{P}/ploy$ by circuit lower bounds). In view of this, we believe that only God knows whether the circuit lower bounds really have a natural proofs barrier to the $\mathcal{P}$ versus $\mathcal{NP}$ problem.
@@
More interestingly, the technique of diagonalization led to some early successes in computational complexity theory. In history, it was invented by Georg Cantor [Can91] to show that there are more real numbers than algebraic numbers, then it was used by Alan Turing to show that there exists a sequence that is not computable (see [Tur37], p. 246), and then refined to prove computational complexity lower bounds by Hartmanis and Stearns [HS65]. A typical theorem in this area is that more time/space buys more computational power [HS65, Coo73, FS07, Pap94, Sip13, AB09]. For instance, there are functions computable in time $n^{2}$, say, which are not computable in time $n$. At present, we assume that the heart of such arguments is the existence of a universal Turing machine, which can simulate every other Turing machine with only a small loss in efficiency.111 At the end of the work, it will become clear that the notion of Enumerable of Turing machines is an important prerequisite for application of diagonalization.
@@
As we mentioned earlier, the diagonalization technique was invented by Cantor [Can91] to show that there are more real numbers than algebraic numbers, and early it demonstrated powers in separating two complexity classes. There is an important doubt: can the diagonalization approach resolve the $\mathcal{P}$ versus $\mathcal{NP}$ problem? Researchers realized in the $1970$s that diagonalization alone may not be able to resolve the $\mathcal{P}$ versus $\mathcal{NP}$ problem; it needs other new techniques besides diagonalization (see e.g., [Sip13, AB09]). But, at the end of the work, we will say that the above point of view is partially correct and leave aside why we think so.
@@
It is interesting that the complexity class $\mathcal{NP}$ has a rich structure under the assumption that $\mathcal{P}$ and $\mathcal{NP}$ differ. In [Lad75], Lander constructed a language that is $\mathcal{NP}$-intermediate by the method of lazy diagonalization under the assumption that $\mathcal{P}\neq\mathcal{NP}$ (there are other problems that are suspected of being $\mathcal{NP}$-intermediate; for example, the graph isomorphism problem is believed to be at least not $\mathcal{NP}$-complete [AK06]). We noted that the lazy diagonalization put the $\mathcal{NP}$-intermediate language accepted by no polynomial-time deterministic Turing machine, and our curiosity here, after reading [Lad75] many times, is why Lander did not diagonalize against all polynomial-time deterministic Turing machines directly to see whether it is possible to give a solution of the $\mathcal{P}$ versus $\mathcal{NP}$ problem. Because it is a widespread belief that $\mathcal{P}$ and $\mathcal{NP}$ are different, we naturally would consider whether we can diagonalize against all of the polynomial-time deterministic Turing machines by a universal nondeterministic Turing machine to produce a language accepted by no polynomial-time deterministic Turing machine but accepted by some nondeterministic Turing machine. Once such a language that is not in $\mathcal{P}$ has been constructed by diagonalization techniques, we can do the second half of the work to see whether this language is in $\mathcal{NP}$ or not.
@@
Before introducing the main results, let us stress that the definitions of almost all of the formal concepts involved in the subsection below will be postponed, and, due to our style of writing, we just suppose that the reader is familiar with these notions. To those who are not familiar with these concepts, we suggest proceeding to read the relevant definitions in Section 2 and in Section 7 first.
1.1. Main Results
@@
We put our aforementioned curiosities, or ideas, into practice in this work. We first enumerate all polynomial-time deterministic Turing machines222 It is much more suitable to say that we first present a way to enumerate all of the polynomial-time deterministic Turing machines. and then diagonalize against all of them by a universal nondeterministic Turing machine. Generally, complexity theorists used universal deterministic Turing machines to diagonalize against a list of deterministic Turing machines (see e.g., [HS65, Pap94, AB09]), or applied universal nondeterministic Turing machines to diagonalize against a list of nondeterministic Turing machines; see e.g., [AB09, For00, FS07, Zak83]. Using a universal nondeterministic Turing machine to diagonalize against a collection of deterministic Turing machines appears less often in the previous literature, except in the author’s recent work [Lin21], and it is maybe a new attempt that may lead to success. As an amazing result, we obtain the following important theorem:
Theorem 1.1.
@@
There exists a language $L_{d}$ that is not accepted by any polynomial-time deterministic Turing machines but is accepted by a nondeterministic Turing machine. Furthermore, this nondeterministic Turing machine runs within time $O(n^{k})$ for any $k\in\mathbb{N}_{1}$. By this, it can be shown that $L_{d}\in\mathcal{NP}$.
From which it immediately follows that:
Corollary 1.2.
@@
In order to contrast with the construction in the proof of Theorem 1.1, we also design a universal deterministic Turing machine $M^{\prime}_{0}$ that accepts a language in $\mathcal{P}$, and this machine also runs within time $O(n^{k})$ for all $k\in\mathbb{N}_{1}$. In contrast, the deterministic universal Turing machine $M^{\prime}_{0}$ does not perform any diagonalization operation. More interestingly, we conjecture that the language accepted by this deterministic universal Turing machine $M^{\prime}_{0}$ cannot be accepted by any deterministic $O(n^{i})$ time-bounded Turing machine for fixed $i\in\mathbb{N}_{1}$. Thus, the following theorem is an interesting by-product of our similar but different construction in proof of Theorem 1.1:
Theorem 1.4.
@@
Now, let us turn to the “Relativization Barrier.”
@@
In $1975$, Baker, Gill, and Solovay [BGS75] presented a proof of that:333 The proof is via $\mathcal{PSPACE}=\mathcal{NPSPACE}$, whose proof is by the fact that the space is reusable but not valid for time, obviously.
@@
What are the implications of the above relativized result for the corresponding unrelativized question? A number of interesting ideas have been proposed, as observed by Ko (see [Ko85]): for example, Baker, Gill, and Solovay [BGS75] suggested that their result implies that ordinary diagonalization techniques are not capable of proving $\mathcal{P}\neq\mathcal{NP}$ (similar perspectives also followed in [Sip13, BC94]). However, Kozen [Koz78] disagreed with this point of view. Hartmanis and Hopcroft [HH76] pointed out the possibility of the axiomatic independence of the question $\mathcal{P}\overset{?}{=}\mathcal{NP}$. Bennett and Gill [BG81] showed that $\mathcal{P}^{B}=\mathcal{R}^{B}$ but $\mathcal{P}^{B}\neq\mathcal{NP}^{B}$ relative to a random oracle $B$. Based on their results, they proposed the random oracle hypothesis: for an “acceptable” relativized statement $S^{A}$, $S^{\emptyset}$ is true if and only if $S^{A}$ is true with probability $1$ when $A$ is random.
@@
To this day, almost all complexity theory experts notice that the proof techniques used to prove $\mathcal{P}\neq EXP$, i.e., diagonalization techniques, would also ‘apply verbatim’ if we added an arbitrary oracle $O$. Thus, for any oracle $O$, we have $\mathcal{P}^{O}\neq EXP^{O}$. However, if we used similar techniques to show that $\mathcal{P}\neq\mathcal{NP}$, then it would also follow that $\mathcal{P}^{O}\neq\mathcal{NP}^{O}$ for all oracle $O$, which contradicts the result of [BGS75] above. This is the so-called “Relativization Barrier,” which almost all complexity theory experts think that any proof technique leading to $\mathcal{P}\neq\mathcal{NP}$ should overcome. Simultaneously, this is also the most notable difficulty on the road to attacking this problem in the direction of $\mathcal{P}\neq\mathcal{NP}$ prior to our work. Based on the original perspectives given in [BGS75], it seems that the motivation of [BGS75] is to prove that $\mathcal{P}\neq\mathcal{NP}$ implies that $\mathcal{P}^{O}\neq\mathcal{NP}^{O}$ for all oracle $O$, via relativization, i.e., a proof technique invariant to adding oracles, but nevertheless we prefer to regard that $\mathcal{P}\neq\mathcal{NP}$ is not necessarily a necessary and sufficient condition for $\mathcal{P}^{O}\neq\mathcal{NP}^{O}$ for all oracle $O$444 For example, in the case where the set of $P^{O}$ is not enumerable, as will be explained below in detail. Likewise, we prefer the point of view that the conclusions [BGS75, Yao85] (i.e., there exists an oracle $A$ such that $\mathcal{P}^{A}\neq\mathcal{NP}^{A}$) imply that $\mathcal{P}^{A}=\mathcal{NP}^{A}$ for all oracle $A$ is not necessarily a necessary and sufficient condition for $\mathcal{P}=\mathcal{NP}$.
@@
On the other hand, the role of relativization in complexity theory is interesting and important and has been a central theme in complexity theory for almost two decades. Following the work [BGS75], much effort was expended to find contradictory relativizations for other open problems in complexity theory. A major part of this effort revolved around obtaining relativized results about the polynomial time hierarchy and its relationship to other classes; see e.g., [HCCRR93]. For example, contradictory relativizations for various unsolved problems regarding the polynomial hierarchy were found (see e.g., [HCCRR93]). These and other results led to a strong belief that problems with contradictory relativization are very hard to solve and are not amenable to current proof techniques, i.e., the solutions of such problems are beyond the current state of mathematics (see e.g., [Hop84, HCCRR93]).
@@
Now let us return to the proof technique of diagonalization once again. Cantor’s diagonal process, also called the diagonalization argument, was published in $1891$ by Georg Cantor [Can91] as a mathematical proof that there are infinite sets that cannot be put into one-to-one correspondence with the infinite set of positive numbers, i.e., $\mathbb{N}_{1}$ defined in the following Section 2. The technique of diagonalization was first used in computability theory in the $1930$s by Turing [Tur37] to show that there exists undecidable language. In computational complexity theory, in their seminal paper [HS65], Hartmanis and Stearns employed the diagonalization proof to give time hierarchy. For more other summarization about it, please consult [For00] for a survey. On the other hand, some complexity theorists or textbooks (see e.g., page $73$ in [AB09]) in computational complexity theory regard that “diagonalization” is any technique that relies solely upon the following properties of Turing machines:
I: